OpenStudy (anonymous):

The sum of the first n consecutive natural numbers is n(n=1) all over 2. For example, 1+2+3+4=10 4(4+1) all over 2. Find all values of n with the property that this sum lies between 150 and 200.

OpenStudy (anonymous):

150*2=300 200*2=400 we have to find two consecutive numbers whose product lies between 300 &400 16*17=272 17*18=306 18*19=342 19*20=380 20*21=420 value of n=17,18,19